# DriveRX: A Vision-Language Reasoning Model for Cross-Task Autonomous Driving

### v1

@@ I.2 Inference Prompt
Method
Size
$\bullet$ Clean
$\bullet$ Brightness
$\bullet$ Dark
$\bullet$ Snow
$\bullet$ Fog
$\bullet$ Rain
$\bullet$ Lens Obstacle
$\bullet$ Water Splash
$\bullet$ Camera Crash
$\bullet$ Frame Lost
$\bullet$ Saturate
$\bullet$ Motion Blur
$\bullet$ Zoom Blur
$\bullet$ Bit Error
$\bullet$ Color Quant
$\bullet$ H.265 Compression
LLaVA-1.5 $7$B $32.40$ $32.48$ $32.95$ $31.95$ $32.43$ $32.30$ $32.88$ $32.18$ $32.93$ $31.63$ $32.50$ $32.43$ $32.93$ $32.48$ $32.18$ $32.63$
LLaVA-NeXT $7$B $32.98$ $33.85$ $20.43$ $11.62$ $16.58$ $27.33$ $18.24$ $32.30$ $26.80$ $20.83$ $23.50$ $34.00$ $17.75$ $24.50$ $18.03$ $26.24$
Qwen2.5-VL $7$B $36.75$ $37.26$ $24.91$ $30.21$ $24.36$ $32.41$ $27.53$ $26.71$ $26.50$ $19.03$ $33.42$ $34.50$ $18.15$ $22.41$ $21.47$ $38.41$
InternVL3 $8$B $36.22$ $39.97$ $35.94$ $42.59$ $37.24$ $41.41$ $41.24$ $39.39$ $33.74$ $26.38$ $35.06$ $39.82$ $41.33$ $31.00$ $40.65$ $42.12$
Qwen2.5VL $72$B $47.68$ $39.71$ $26.35$ $38.59$ $42.56$ $45.03$ $38.41$ $39.38$ $28.21$ $19.38$ $44.18$ $39.85$ $36.76$ $32.35$ $29.55$ $49.47$
InternVL3 $78$B $47.33$ $44.82$ $38.35$ $45.29$ $43.50$ $40.09$ $46.09$ $41.00$ $32.09$ $24.53$ $40.68$ $47.30$ $39.44$ $36.61$ $41.30$ $45.97$
MM-Eureka $7$B $43.33$ $29.21$ $17.59$ $31.97$ $37.97$ $41.44$ $36.44$ $32.21$ $30.27$ $22.71$ $41.73$ $34.68$ $19.56$ $26.39$ $27.21$ $41.21$
R1-Onevision $7$B $36.89$ $30.24$ $23.71$ $31.56$ $37.36$ $34.32$ $36.12$ $25.06$ $30.18$ $34.74$ $26.82$ $24.76$ $34.03$ $23.71$ $29.56$ $29.59$
DriveLM $7$B $22.38$ $20.78$ $25.30$ $18.98$ $24.43$ $25.95$ $22.03$ $21.03$ $21.95$ $16.28$ $19.38$ $22.98$ $20.93$ $19.90$ $16.25$ $26.48$
Dolphin $7$B $6.50$ $10.18$ $11.08$ $10.70$ $9.53$ $10.58$ $9.93$ $9.80$ $10.08$ $9.95$ $11.20$ $9.85$ $10.10$ $8.80$ $10.00$ $11.10$
Align-DS-V $8$B $39.8$ $39.65$ $31.00$ $41.50$ $33.24$ $42.50$ $35.15$ $36.85$ $27.41$ $38.32$ $25.21$ $28.68$ $30.44$ $35.91$
DriveRX $8$B $42.79$ $34.71$ $36.09$ $39.29$ $43.74$ $39.79$ $27.56$ $38.85$ $33.68$ $26.74$ $32.18$ $33.29$ $30.53$ $30.65$ $25.59$ $38.50$
@@
LLaVA-1.5 $7$B $14.03$ $13.53$ $13.31$ $13.31$ $13.61$ $13.75$ $13.91$ $13.48$ $13.95$ $12.90$ $12.98$ $13.35$ $13.05$ $13.36$ $12.83$ $14.28$
LLaVA-NeXT $7$B $15.33$ $15.49$ $14.95$ $16.61$ $15.62$ $16.05$ $15.66$ $15.66$ $15.19$ $16.04$ $15.16$ $16.06$ $17.92$ $15.27$ $15.10$ $15.86$
Qwen2.5-VL $7$B $27.09$ $27.76$ $23.14$ $27.05$ $28.05$ $25.90$ $28.24$ $28.29$ $28.48$ $25.81$ $28.62$ $25.90$ $25.86$ $25.38$ $23.62$ $25.38$
InternVL3 $8$B $30.21$ $30.38$ $31.29$ $32.33$ $28.52$ $29.57$ $28.95$ $31.33$ $29.95$ $25.29$ $30.76$ $29.76$ $30.67$ $28.71$ $29.90$ $30.33$
Qwen2.5VL $72$B $27.43$ $29.29$ $26.76$ $23.86$ $25.05$ $29.38$ $28.14$ $28.43$ $29.38$ $26.14$ $27.86$ $28.71$ $28.10$ $27.52$ $29.48$ $29.71$
InternVL3 $78$B $31.53$ $32.33$ $31.95$ $32.05$ $29.62$ $30.38$ $33.35$ $28.76$ $32.86$ $28.76$ $29.14$ $28.90$ $30.71$ $30.76$ $30.90$ $31.10$
MM-Eureka $7$B $24.46$ $26.48$ $25.38$ $26.29$ $26.33$ $26.62$ $27.71$ $27.14$ $26.10$ $26.00$ $26.19$ $27.14$ $26.52$ $25.67$ $26.81$ $26.62$
R1-Onevision $7$B $20.39$ $21.39$ $19.20$ $19.52$ $19.70$ $20.60$ $21.50$ $21.89$ $22.83$ $20.95$ $20.84$ $21.68$ $17.93$ $18.67$ $22.00$ $21.67$
DriveLM $7$B $11.32$ $11.30$ $10.03$ $11.21$ $9.71$ $10.22$ $10.71$ $11.01$ $11.14$ $10.13$ $9.38$ $10.97$ $9.03$ $11.39$ $10.50$ $10.73$
Dolphin $7$B $12.68$ $12.07$ $10.34$ $12.37$ $11.46$ $11.73$ $12.33$ $11.04$ $12.34$ $11.39$ $11.47$ $11.34$ $10.17$ $11.40$ $11.35$ $11.65$
Align-DS-V $8$B $23.03$ $39.65$ $31.00$ $41.50$ $33.24$ $42.50$ $35.15$ $36.85$ $27.41$ $38.32$ $25.21$ $28.68$ $30.44$ $35.91$ $36.32$ $25.71$
DriveRX $8$B $22.97$ $24.19$ $26.33$ $26.33$ $27.52$ $25.86$ $25.38$ $25.14$ $26.10$ $26.81$ $25.38$ $27.05$ $24.67$ $23.71$ $25.76$ $25.62$
@@
LLaVA-1.5 $7$B $22.02$ $24.79$ $20.95$ $15.97$ $15.30$ $18.98$ $16.28$ $24.16$ $6.11$ $13.90$ $20.30$ $25.20$ $10.56$ $11.10$ $15.61$ $23.92$
LLaVA-NeXT $7$B $35.07$ $37.15$ $35.31$ $37.59$ $37.62$ $35.44$ $37.00$ $35.87$ $36.25$ $30.10$ $40.56$ $34.66$ $39.36$ $31.74$ $34.07$ $35.66$
Qwen2.5-VL $7$B $42.26$ $50.75$ $36.58$ $47.25$ $46.00$ $47.58$ $54.50$ $53.67$ $45.00$ $46.17$ $44.83$ $48.08$ $43.00$ $48.00$ $43.25$ $52.25$
InternVL3 $8$B $41.21$ $50.17$ $43.08$ $51.50$ $54.50$ $45.75$ $47.67$ $48.33$ $39.17$ $45.17$ $46.92$ $46.33$ $48.33$ $46.58$ $50.42$ $54.17$
Qwen2.5VL $72$B $54.89$ $56.00$ $46.58$ $38.83$ $44.00$ $51.50$ $57.33$ $54.42$ $47.17$ $52.83$ $50.58$ $50.50$ $55.75$ $50.42$ $51.17$ $49.92$
InternVL3 $78$B $50.93$ $55.50$ $44.92$ $51.50$ $59.00$ $53.92$ $51.67$ $50.33$ $46.92$ $42.83$ $49.08$ $47.75$ $51.42$ $45.08$ $47.75$ $48.00$
MM-Eureka $7$B $40.31$ $44.42$ $37.58$ $36.00$ $36.75$ $32.42$ $42.91$ $45.42$ $45.17$ $53.75$ $45.33$ $43.92$ $39.75$ $39.17$ $38.75$ $44.50$
R1-Onevision $7$B $51.22$ $44.40$ $33.25$ $47.92$ $58.17$ $47.75$ $46.33$ $47.42$ $48.80$ $46.25$ $42.33$ $46.27$ $47.45$ $52.08$ $45.10$ $50.18$
DriveLM $7$B $44.33$ $46.82$ $43.90$ $42.33$ $35.84$ $44.13$ $44.00$ $42.59$ $46.25$ $33.56$ $29.69$ $42.15$ $19.00$ $38.20$ $44.33$ $42.87$
Dolphin $7$B $32.66$ $29.85$ $32.31$ $24.64$ $29.92$ $31.38$ $33.41$ $31.79$ $29.05$ $30.93$ $30.49$ $31.59$ $26.38$ $30.13$ $25.64$ $30.62$
Align-DS-V $8$B $41.51$ $41.50$ $35.83$ $38.75$ $32.67$ $39.42$ $42.50$ $38.33$ $41.92$ $32.83$ $42.92$ $44.42$ $35.25$ $40.67$ $35.42$ $35.00$
DriveRX $8$B $52.20$ $51.42$ $41.75$ $48.83$ $48.92$ $47.17$ $49.25$ $51.92$ $47.92$ $38.67$ $55.42$ $48.17$ $53.25$ $48.50$ $39.50$ $53.75$
@@
LLaVA-1.5 $7$B $29.15$ $31.52$ $31.49$ $32.58$ $32.42$ $31.00$ $29.81$ $31.72$ $33.70$ $35.95$ $29.93$ $31.20$ $30.65$ $30.05$ $30.00$ $30.61$
LLaVA-NeXT $7$B $45.27$ $45.64$ $44.54$ $43.55$ $44.17$ $45.08$ $44.62$ $44.21$ $45.69$ $44.51$ $41.17$ $45.30$ $44.57$ $43.67$ $43.57$ $45.13$
Qwen2.5-VL $7$B $55.70$ $62.34$ $61.15$ $61.76$ $56.91$ $54.27$ $57.51$ $57.48$ $56.37$ $48.50$ $65.56$ $60.16$ $56.63$ $58.27$ $54.06$ $58.60$
InternVL3 $8$B $70.97$ $54.07$ $51.07$ $50.00$ $53.82$ $55.96$ $49.00$ $49.57$ $54.77$ $52.09$ $50.56$ $49.07$ $48.49$ $51.88$ $49.99$ $52.94$
Qwen2.5VL $72$B $77.18$ $80.95$ $69.67$ $75.22$ $76.24$ $77.79$ $78.12$ $78.40$ $69.98$ $53.11$ $74.74$ $80.84$ $75.13$ $73.33$ $68.98$ $77.67$
InternVL3 $78$B $82.24$ $79.30$ $73.96$ $76.63$ $79.51$ $80.70$ $82.28$ $81.33$ $77.95$ $69.71$ $77.17$ $82.17$ $79.67$ $78.41$ $77.48$ $79.90$
MM-Eureka $7$B $65.38$ $63.35$ $58.75$ $60.85$ $59.21$ $66.11$ $57.41$ $62.58$ $59.90$ $49.80$ $60.99$ $62.35$ $58.98$ $57.99$ $52.95$ $71.53$
R1-Onevision $7$B $55.12$ $53.03$ $55.92$ $58.28$ $58.74$ $57.04$ $57.44$ $56.36$ $57.13$ $60.25$ $53.11$ $49.87$ $55.22$ $51.55$ $50.57$ $58.60$
DriveLM $7$B $68.71$ $67.25$ $67.52$ $65.72$ $63.08$ $69.60$ $69.04$ $67.97$ $67.85$ $66.47$ $66.25$ $67.93$ $70.17$ $68.46$ $68.30$ $68.59$
Dolphin $7$B $52.91$ $51.85$ $55.39$ $53.09$ $54.78$ $53.92$ $51.79$ $53.57$ $55.73$ $57.81$ $55.78$ $51.42$ $50.95$ $53.35$ $54.89$ $52.17$
Align-DS-V $8$B $51.16$ $61.88$ $56.82$ $53.65$ $58.77$ $59.20$ $58.10$ $60.98$ $58.71$ $57.55$ $51.16$ $59.95$ $55.45$ $60.63$ $53.04$ $60.82$
DriveRX $8$B $78.63$ $77.89$ $74.99$ $74.54$ $74.33$ $75.29$ $75.44$ $74.71$ $75.96$ $78.18$ $74.87$ $74.18$ $76.48$ $75.59$ $73.67$ $76.80$
@@
LLaVA-1.5 $7$B $13.60$ $12.79$ $12.83$ $15.57$ $12.63$ $14.06$ $13.99$ $12.79$ $14.68$ $13.65$ $13.12$ $13.55$ $13.83$ $13.98$ $13.44$ $13.48$
LLaVA-NeXT $7$B $48.16$ $48.84$ $38.82$ $15.90$ $39.13$ $47.07$ $20.72$ $47.02$ $48.20$ $36.67$ $39.69$ $47.36$ $39.60$ $46.99$ $28.13$ $47.55$
Qwen2.5-VL $7$B $47.40$ $50.59$ $30.63$ $46.11$ $45.26$ $46.30$ $36.26$ $37.85$ $54.56$ $49.41$ $44.81$ $39.56$ $36.81$ $33.59$ $48.89$ $41.67$
InternVL3 $8$B $51.47$ $40.22$ $62.26$ $54.07$ $57.48$ $48.59$ $54.96$ $53.93$ $48.78$ $49.70$ $43.04$ $51.89$ $59.70$ $41.70$ $51.30$ $60.15$
Qwen2.5VL $72$B $55.57$ $45.04$ $58.67$ $59.48$ $57.56$ $51.56$ $49.67$ $57.63$ $57.93$ $49.37$ $53.52$ $49.44$ $35.19$ $41.59$ $56.37$ $51.70$
InternVL3 $78$B $50.73$ $47.70$ $56.04$ $42.41$ $55.78$ $53.07$ $50.67$ $51.41$ $59.30$ $51.19$ $68.19$ $42.85$ $37.04$ $54.07$ $58.59$ $52.15$
MM-Eureka $7$B $49.4$ $47.58$ $54.05$ $36.73$ $39.04$ $59.79$ $46.11$ $51.45$ $50.86$ $61.00$ $42.00$ $56.93$ $30.61$ $46.48$ $34.47$ $56.80$
R1-Onevision $7$B $41.57$ $39.35$ $57.17$ $38.31$ $31.74$ $48.33$ $45.15$ $51.23$ $44.00$ $37.07$ $37.92$ $30.81$ $32.04$ $35.85$ $32.26$ $36.93$
DriveLM $7$B $42.78$ $47.18$ $36.30$ $40.70$ $39.18$ $40.93$ $43.30$ $40.98$ $39.95$ $38.23$ $40.08$ $45.68$ $38.88$ $41.10$ $33.50$ $39.65$
Dolphin $7$B $8.81$ $7.17$ $9.54$ $9.02$ $6.48$ $8.05$ $7.95$ $7.10$ $9.29$ $8.94$ $8.02$ $8.02$ $9.42$ $8.37$ $10.07$ $6.32$
Align-DS-V $8$B $50.53$ $50.26$ $36.07$ $38.04$ $38.67$ $44.78$ $28.52$ $54.41$ $48.19$ $34.30$ $30.96$ $36.22$ $35.56$ $52.33$ $34.52$ $50.63$
DriveRX $8$B $62.02$ $55.67$ $61.04$ $54.07$ $57.15$ $60.35$ $56.22$ $54.70$ $58.63$ $51.19$ $46.85$ $57.30$ $41.15$ $45.88$ $69.04$ $63.48$
@@
Appendix K Implementation Details
K.1 Hardware Configuration
@@
K.2 Hyperparameter Settings
During GRPO [40] training, we adopt a prompt batch size of 128 is utilized, with 8 rollouts generated per prompt. The maximum rollout length is capped at 4,096 tokens. A default sampling temperature of 1.0 is set, and a KL loss coefficient of 1e-2 is employed. The training process consists of 5 epochs.
@@
To ensure the reliability of the evaluation, each model is tested three times, with the average value taken as the final result. For evaluation purposes, all models are adjusted to a temperature of 0.2. The vLLM[57] framework is employed for inference, with the maximum length restricted to 4096 tokens.
Appendix L Comparison with DriveLMM-o1 Using DriveBench Evaluation
DriveLMM-o1[58] is a recent work closely related to ours, also targeting multi-task visual-language reasoning in autonomous driving. It constructs a step-by-step reasoning dataset covering perception, prediction, and planning tasks, and fine-tunes an InternVL2.5-8B[59] model using LoRA to improve logical coherence and interpretability in decision making.

### v2

@@ Appendix I Alignment Between Human and Model Rewards
LLaVA-1.5 77B $32.40$ $32.48$ $32.95$ $31.95$ $32.43$ $32.30$ $32.88$ $32.18$ $32.93$ $31.63$ $32.50$ $32.43$ $32.93$ $32.48$ $32.18$ $32.63$
LLaVA-NeXT 77B $32.98$ $33.85$ $20.43$ $11.62$ $16.58$ $27.33$ $18.24$ $32.30$ $26.80$ $20.83$ $23.50$ $34.00$ $17.75$ $24.50$ $18.03$ $26.24$
Qwen2.5-VL 77B $36.75$ $37.26$ $24.91$ $30.21$ $24.36$ $32.41$ $27.53$ $26.71$ $26.50$ $19.03$ $33.42$ $34.50$ $18.15$ $22.41$ $21.47$ $38.41$
InternVL3 88B $36.22$ $39.97$ $35.94$ $42.59$ $37.24$ $41.41$ $41.24$ $39.39$ $33.74$ $26.38$ $35.06$ $39.82$ $41.33$ $31.00$ $40.65$ $42.12$
Qwen2.5VL 7272B $47.68$ $39.71$ $26.35$ $38.59$ $42.56$ $45.03$ $38.41$ $39.38$ $28.21$ $19.38$ $44.18$ $39.85$ $36.76$ $32.35$ $29.55$ $49.47$
InternVL3 7878B $47.33$ $44.82$ $38.35$ $45.29$ $43.50$ $40.09$ $46.09$ $41.00$ $32.09$ $24.53$ $40.68$ $47.30$ $39.44$ $36.61$ $41.30$ $45.97$
$M$-Eureka 77B $43.33$ $29.21$ $17.59$ $31.97$ $37.97$ $41.44$ $36.44$ $32.21$ $30.27$ $22.71$ $41.73$ $34.68$ $19.56$ $26.39$ $27.21$ $41.21$
R1-Onevision 77B $36.89$ $30.24$ $23.71$ $31.56$ $37.36$ $34.32$ $36.12$ $25.06$ $30.18$ $34.74$ $26.82$ $24.76$ $34.03$ $23.71$ $29.56$ $29.59$
DriveLM 77B $22.38$ $20.78$ $25.30$ $18.98$ $24.43$ $25.95$ $22.03$ $21.03$ $21.95$ $16.28$ $19.38$ $22.98$ $20.93$ $19.90$ $16.25$ $26.48$
Dolphin 77B $6.50$ $10.18$ $11.08$ $10.70$ $9.53$ $10.58$ $9.93$ $9.80$ $10.08$ $9.95$ $11.20$ $9.85$ $10.10$ $8.80$ $10.00$ $11.10$
Align-DS-V 88B $39.8$ $39.65$ $31.00$ $41.50$ $33.24$ $42.50$ $35.15$ $36.85$ $27.41$ $38.32$ $25.21$ $28.68$ $30.44$ $35.91$
DriveRX 88B $42.79$ $34.71$ $36.09$ $39.29$ $43.74$ $39.79$ $27.56$ $38.85$ $33.68$ $26.74$ $32.18$ $33.29$ $30.53$ $30.65$ $25.59$ $38.50$
@@
LLaVA-1.5 77B $14.03$ $13.53$ $13.31$ $13.31$ $13.61$ $13.75$ $13.91$ $13.48$ $13.95$ $12.90$ $12.98$ $13.35$ $13.05$ $13.36$ $12.83$ $14.28$
LLaVA-NeXT 77B $15.33$ $15.49$ $14.95$ $16.61$ $15.62$ $16.05$ $15.66$ $15.66$ $15.19$ $16.04$ $15.16$ $16.06$ $17.92$ $15.27$ $15.10$ $15.86$
Qwen2.5-VL 77B $27.09$ $27.76$ $23.14$ $27.05$ $28.05$ $25.90$ $28.24$ $28.29$ $28.48$ $25.81$ $28.62$ $25.90$ $25.86$ $25.38$ $23.62$ $25.38$
InternVL3 88B $30.21$ $30.38$ $31.29$ $32.33$ $28.52$ $29.57$ $28.95$ $31.33$ $29.95$ $25.29$ $30.76$ $29.76$ $30.67$ $28.71$ $29.90$ $30.33$
Qwen2.5VL 7272B $27.43$ $29.29$ $26.76$ $23.86$ $25.05$ $29.38$ $28.14$ $28.43$ $29.38$ $26.14$ $27.86$ $28.71$ $28.10$ $27.52$ $29.48$ $29.71$
InternVL3 7878B $31.53$ $32.33$ $31.95$ $32.05$ $29.62$ $30.38$ $33.35$ $28.76$ $32.86$ $28.76$ $29.14$ $28.90$ $30.71$ $30.76$ $30.90$ $31.10$
$M$-Eureka 77B $24.46$ $26.48$ $25.38$ $26.29$ $26.33$ $26.62$ $27.71$ $27.14$ $26.10$ $26.00$ $26.19$ $27.14$ $26.52$ $25.67$ $26.81$ $26.62$
R1-Onevision 77B $20.39$ $21.39$ $19.20$ $19.52$ $19.70$ $20.60$ $21.50$ $21.89$ $22.83$ $20.95$ $20.84$ $21.68$ $17.93$ $18.67$ $22.00$ $21.67$
DriveLM 77B $11.32$ $11.30$ $10.03$ $11.21$ $9.71$ $10.22$ $10.71$ $11.01$ $11.14$ $10.13$ $9.38$ $10.97$ $9.03$ $11.39$ $10.50$ $10.73$
Dolphin 77B $12.68$ $12.07$ $10.34$ $12.37$ $11.46$ $11.73$ $12.33$ $11.04$ $12.34$ $11.39$ $11.47$ $11.34$ $10.17$ $11.40$ $11.35$ $11.65$
Align-DS-V 88B $23.03$ $39.65$ $31.00$ $41.50$ $33.24$ $42.50$ $35.15$ $36.85$ $27.41$ $38.32$ $25.21$ $28.68$ $30.44$ $35.91$ $36.32$ $25.71$
DriveRX 88B $22.97$ $24.19$ $26.33$ $26.33$ $27.52$ $25.86$ $25.38$ $25.14$ $26.10$ $26.81$ $25.38$ $27.05$ $24.67$ $23.71$ $25.76$ $25.62$
@@
LLaVA-1.5 77B $22.02$ $24.79$ $20.95$ $15.97$ $15.30$ $18.98$ $16.28$ $24.16$ $6.11$ $13.90$ $20.30$ $25.20$ $10.56$ $11.10$ $15.61$ $23.92$
LLaVA-NeXT 77B $35.07$ $37.15$ $35.31$ $37.59$ $37.62$ $35.44$ $37.00$ $35.87$ $36.25$ $30.10$ $40.56$ $34.66$ $39.36$ $31.74$ $34.07$ $35.66$
Qwen2.5-VL 77B $42.26$ $50.75$ $36.58$ $47.25$ $46.00$ $47.58$ $54.50$ $53.67$ $45.00$ $46.17$ $44.83$ $48.08$ $43.00$ $48.00$ $43.25$ $52.25$
InternVL3 88B $41.21$ $50.17$ $43.08$ $51.50$ $54.50$ $45.75$ $47.67$ $48.33$ $39.17$ $45.17$ $46.92$ $46.33$ $48.33$ $46.58$ $50.42$ $54.17$
Qwen2.5VL 7272B $54.89$ $56.00$ $46.58$ $38.83$ $44.00$ $51.50$ $57.33$ $54.42$ $47.17$ $52.83$ $50.58$ $50.50$ $55.75$ $50.42$ $51.17$ $49.92$
InternVL3 7878B $50.93$ $55.50$ $44.92$ $51.50$ $59.00$ $53.92$ $51.67$ $50.33$ $46.92$ $42.83$ $49.08$ $47.75$ $51.42$ $45.08$ $47.75$ $48.00$
$M$-Eureka 77B $40.31$ $44.42$ $37.58$ $36.00$ $36.75$ $32.42$ $42.91$ $45.42$ $45.17$ $53.75$ $45.33$ $43.92$ $39.75$ $39.17$ $38.75$ $44.50$
R1-Onevision 77B $51.22$ $44.40$ $33.25$ $47.92$ $58.17$ $47.75$ $46.33$ $47.42$ $48.80$ $46.25$ $42.33$ $46.27$ $47.45$ $52.08$ $45.10$ $50.18$
DriveLM 77B $44.33$ $46.82$ $43.90$ $42.33$ $35.84$ $44.13$ $44.00$ $42.59$ $46.25$ $33.56$ $29.69$ $42.15$ $19.00$ $38.20$ $44.33$ $42.87$
Dolphin 77B $32.66$ $29.85$ $32.31$ $24.64$ $29.92$ $31.38$ $33.41$ $31.79$ $29.05$ $30.93$ $30.49$ $31.59$ $26.38$ $30.13$ $25.64$ $30.62$
Align-DS-V 88B $41.51$ $41.50$ $35.83$ $38.75$ $32.67$ $39.42$ $42.50$ $38.33$ $41.92$ $32.83$ $42.92$ $44.42$ $35.25$ $40.67$ $35.42$ $35.00$
DriveRX 88B $52.20$ $51.42$ $41.75$ $48.83$ $48.92$ $47.17$ $49.25$ $51.92$ $47.92$ $38.67$ $55.42$ $48.17$ $53.25$ $48.50$ $39.50$ $53.75$
@@
LLaVA-1.5 77B $29.15$ $31.52$ $31.49$ $32.58$ $32.42$ $31.00$ $29.81$ $31.72$ $33.70$ $35.95$ $29.93$ $31.20$ $30.65$ $30.05$ $30.00$ $30.61$
LLaVA-NeXT 77B $45.27$ $45.64$ $44.54$ $43.55$ $44.17$ $45.08$ $44.62$ $44.21$ $45.69$ $44.51$ $41.17$ $45.30$ $44.57$ $43.67$ $43.57$ $45.13$
Qwen2.5-VL 77B $55.70$ $62.34$ $61.15$ $61.76$ $56.91$ $54.27$ $57.51$ $57.48$ $56.37$ $48.50$ $65.56$ $60.16$ $56.63$ $58.27$ $54.06$ $58.60$
InternVL3 88B $70.97$ $54.07$ $51.07$ $50.00$ $53.82$ $55.96$ $49.00$ $49.57$ $54.77$ $52.09$ $50.56$ $49.07$ $48.49$ $51.88$ $49.99$ $52.94$
Qwen2.5VL 7272B $77.18$ $80.95$ $69.67$ $75.22$ $76.24$ $77.79$ $78.12$ $78.40$ $69.98$ $53.11$ $74.74$ $80.84$ $75.13$ $73.33$ $68.98$ $77.67$
InternVL3 7878B $82.24$ $79.30$ $73.96$ $76.63$ $79.51$ $80.70$ $82.28$ $81.33$ $77.95$ $69.71$ $77.17$ $82.17$ $79.67$ $78.41$ $77.48$ $79.90$
$M$-Eureka 77B $65.38$ $63.35$ $58.75$ $60.85$ $59.21$ $66.11$ $57.41$ $62.58$ $59.90$ $49.80$ $60.99$ $62.35$ $58.98$ $57.99$ $52.95$ $71.53$
R1-Onevision 77B $55.12$ $53.03$ $55.92$ $58.28$ $58.74$ $57.04$ $57.44$ $56.36$ $57.13$ $60.25$ $53.11$ $49.87$ $55.22$ $51.55$ $50.57$ $58.60$
DriveLM 77B $68.71$ $67.25$ $67.52$ $65.72$ $63.08$ $69.60$ $69.04$ $67.97$ $67.85$ $66.47$ $66.25$ $67.93$ $70.17$ $68.46$ $68.30$ $68.59$
Dolphin 77B $52.91$ $51.85$ $55.39$ $53.09$ $54.78$ $53.92$ $51.79$ $53.57$ $55.73$ $57.81$ $55.78$ $51.42$ $50.95$ $53.35$ $54.89$ $52.17$
Align-DS-V 88B $51.16$ $61.88$ $56.82$ $53.65$ $58.77$ $59.20$ $58.10$ $60.98$ $58.71$ $57.55$ $51.16$ $59.95$ $55.45$ $60.63$ $53.04$ $60.82$
DriveRX 88B $78.63$ $77.89$ $74.99$ $74.54$ $74.33$ $75.29$ $75.44$ $74.71$ $75.96$ $78.18$ $74.87$ $74.18$ $76.48$ $75.59$ $73.67$ $76.80$
@@
Method
Size
$\bullet$ Clean
$\bullet$ Brightness
$\bullet$ Dark
$\bullet$ Snow
$\bullet$ Fog
$\bullet$ Rain
$\bullet$ Lens Obstacle
$\bullet$ Water Splash
$\bullet$ Camera Crash
$\bullet$ Frame Lost
$\bullet$ Saturate
$\bullet$ Motion Blur
$\bullet$ Zoom Blur
$\bullet$ Bit Error
$\bullet$ Color Quant
$\bullet$ H.265 Compression
LLaVA-1.5 77B $13.60$ $12.79$ $12.83$ $15.57$ $12.63$ $14.06$ $13.99$ $12.79$ $14.68$ $13.65$ $13.12$ $13.55$ $13.83$ $13.98$ $13.44$ $13.48$
LLaVA-NeXT 77B $48.16$ $48.84$ $38.82$ $15.90$ $39.13$ $47.07$ $20.72$ $47.02$ $48.20$ $36.67$ $39.69$ $47.36$ $39.60$ $46.99$ $28.13$ $47.55$
Qwen2.5-VL 77B $47.40$ $50.59$ $30.63$ $46.11$ $45.26$ $46.30$ $36.26$ $37.85$ $54.56$ $49.41$ $44.81$ $39.56$ $36.81$ $33.59$ $48.89$ $41.67$
InternVL3 88B $51.47$ $40.22$ $62.26$ $54.07$ $57.48$ $48.59$ $54.96$ $53.93$ $48.78$ $49.70$ $43.04$ $51.89$ $59.70$ $41.70$ $51.30$ $60.15$
Qwen2.5VL 7272B $55.57$ $45.04$ $58.67$ $59.48$ $57.56$ $51.56$ $49.67$ $57.63$ $57.93$ $49.37$ $53.52$ $49.44$ $35.19$ $41.59$ $56.37$ $51.70$
InternVL3 7878B $50.73$ $47.70$ $56.04$ $42.41$ $55.78$ $53.07$ $50.67$ $51.41$ $59.30$ $51.19$ $68.19$ $42.85$ $37.04$ $54.07$ $58.59$ $52.15$
$M$-Eureka 77B $49.4$ $47.58$ $54.05$ $36.73$ $39.04$ $59.79$ $46.11$ $51.45$ $50.86$ $61.00$ $42.00$ $56.93$ $30.61$ $46.48$ $34.47$ $56.80$
R1-Onevision 77B $41.57$ $39.35$ $57.17$ $38.31$ $31.74$ $48.33$ $45.15$ $51.23$ $44.00$ $37.07$ $37.92$ $30.81$ $32.04$ $35.85$ $32.26$ $36.93$
DriveLM 77B $42.78$ $47.18$ $36.30$ $40.70$ $39.18$ $40.93$ $43.30$ $40.98$ $39.95$ $38.23$ $40.08$ $45.68$ $38.88$ $41.10$ $33.50$ $39.65$
Dolphin 77B $8.81$ $7.17$ $9.54$ $9.02$ $6.48$ $8.05$ $7.95$ $7.10$ $9.29$ $8.94$ $8.02$ $8.02$ $9.42$ $8.37$ $10.07$ $6.32$
Align-DS-V 88B $50.53$ $50.26$ $36.07$ $38.04$ $38.67$ $44.78$ $28.52$ $54.41$ $48.19$ $34.30$ $30.96$ $36.22$ $35.56$ $52.33$ $34.52$ $50.63$
DriveRX 88B $62.02$ $55.67$ $61.04$ $54.07$ $57.15$ $60.35$ $56.22$ $54.70$ $58.63$ $51.19$ $46.85$ $57.30$ $41.15$ $45.88$ $69.04$ $63.48$
Appendix O Implementation Details
O.1 Hardware Configuration
@@
O.2 Hyperparameter Settings
During GRPO (Shao et al., 2024) training, we adopt a prompt batch size of 128 is utilized, with 8 rollouts generated per prompt. The maximum rollout length is capped at 4,096 tokens. A default sampling temperature of 1.0 is set, and a KL loss coefficient of 1e-2 is employed. The training process consists of 5 epochs.
@@
To ensure the reliability of the evaluation, each model is tested three times, with the average value taken as the final result. For evaluation purposes, all models are adjusted to a temperature of 0.2. The vLLM(Kwon et al., 2023) framework is employed for inference, with the maximum length restricted to 4096 tokens.
InternVL3 (Zhu et al., 2025) 88B $33.21$ $41.21$ $70.97$ $51.47$
DriveLMM-o1 (Ishaq et al., 2025) 88B $34.31$ $43.95$ $73.7$ $47.62$
DriveRX 88B $32.88$ $52.20$ $78.63$ $62.02$
O.3 Latency of LLM-Based Reward Scoring
During the GRPO training process, we deployed Qwen2.5-72B-Instruct on 8×A100 using the vLLM framework for LLM-based reward scoring, achieving a throughput of 1,024 samples in 5 minutes (0.29 seconds per sample with batch size 128).
Appendix P Comparison with recent work Using DriveBench Evaluation
DriveLMM-o1(Ishaq et al., 2025) is a recent work closely related to ours, also targeting multi-task visual-language reasoning in autonomous driving. It constructs a step-by-step reasoning dataset covering perception, prediction, and planning tasks, and fine-tunes an InternVL2.5-8B(Chen et al., 2024b) model using LoRA to improve logical coherence and interpretability in decision making.